# MD 20742, USA

### v2

@@ 3.1. Splitting the sum
3.3. Controlling small values of $X$ for $21\leq s<$ 2
Here we prove (3.3) in the case $21\leq s<$ 2 .
Given $y\in\mathbb{S}$, let $C_{n}(y)$ be the continuity component of $\phi 1_{\phi\leq N^{1/s}(\operatorname{ln}N)^{-D}}\circ f^{k},$ $k\leq n$ containing $y$ and $f^{n}|_{C_{n}(y)}$ be a diffeomorphism. Let $I_{n}(y)=f^{n}C_{n}(y).$ We say that $C_{n}(y)$ is short if the length of $I_{n}(y)$ is less than $N^{-1}(\operatorname{ln}N)^{-100}$ and we say that it is long otherwise. Fix $\varepsilon>0,$ we say that $C_{n}$ is extremal if there exists $z\in C_{n}$ such that $|{\bar{S}}_{n}(z)|\geq{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}$ and for each $k<n,$ $|{\bar{S}}_{k}(z)|<{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}$ on $C_{n}.$ Note that different extremal components are disjoint. In addition,
(3.13) $|{\bar{S}}_{n}(y)|\geq\frac{3}{4}{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}\quad\text{for all }y\text{ in }C_{n}.$
@@ 3.3. Controlling small values of $X$ for $21\leq s<$ 2
$\left|{\bar{S}}_{n}(y)-{\bar{S}}_{n}(z)\right|\leq\sum_{k=1}^{n-K\operatorname{ln}N}|{\bar{X}}_{k}(y)-{\bar{X}}_{k}(z)|+|{\bar{S}}_{K\operatorname{ln}N}(f^{n-K\operatorname{ln}N}y)|+|{\bar{S}}_{K\operatorname{ln}N}(f^{n-K\operatorname{ln}N}z)|,$
then the last two terms are smaller than $2KN^{1/s}(\operatorname{ln}N)^{1-D}$ due to the cutoff in the definition of ${\bar{S}}_{n}.$ To estimate the first term, we note that since $d(f^{n}y,f^{n}z)\leq 1$ and $f$ is expanding then for $k\leq n-K\operatorname{ln}N,$ $d(f^{k}y,f^{k}z)\leq\theta^{K\operatorname{ln}N}.$ Since $\phi{\mathbf{1}}_{\{\phi<N^{1/s}(\operatorname{ln}N)^{-D}\}}$ is Lipschitz on its continuity components with Lipschitz constant less than $CN^{1+1/s}(\operatorname{ln}N)^{-(1+s)D}$, it follows that if $K$ is sufficiently large then $|{\bar{X}}_{k}(y)-{\bar{X}}_{k}(z)|\leq N^{-100}$, proving (3.13).
@@
If $D$ is large enough then $\mu({\mathcal{E}}_{N})\leq{C}{(\operatorname{ln}N)^{-2}}.$
Lemma 3.2 implies (3.3). In fact, combining this lemma with Borel-Cantelli Lemma we conclude that with probability $1$, for large $k$, $|{\bar{S}}_{n}|<{\varepsilon}N_{k}^{1/s}(\operatorname{ln}N_{k})^{\alpha}$ for $n\leq N_{k},$ $N_{k}=2^{k}.$ Since ${\varepsilon}$ is arbitrary, (3.3) follows.
@@
To prove (3.15) we use Chebyshev inequality. By (2.2) we have
@@
Since $D(2-s)+2\alpha>3$ due to (3.2), (3.15) follows by Chebyshev inequality.
The proof of (3.14) follows similar ideas. Namely, in view of (3.13), it suffices to show that there exists $C<1$ such that
(3.19) $\mu\left(y\in C_{n}:\left|\sum_{k=n+1}^{N}{\bar{X}}_{k}(y)\right|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq C\mu(C_{n}).$
@@
(3.20) $\mu\left(z\in I_{n}:|{\bar{S}}_{N-n}(z)|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq C\mu(I_{n}).$
@@
Combining (3.21) and (3.22) we obtain (3.14). $\square$
3.4. Controlling intermediate values of $X$ for $21\leq s<$ 2
Here we prove (3.4) in case $21\leq s<$ 2 .
@@ 3.4. Controlling intermediate values of $X$ for $21\leq s<$ 2
[…]
@@
Equation (3.24) follows from the Markov inequality and the following estimate.
@@
Suppose $x$ is Diophantine. For each $m,$ we have
@@
Note that by (2.2)
(3.30) ${\mathbb{E}}\left|\bar{\bar{X}}_{0}\right|^{l}\leq\left\{\begin{array}[]{ll}CN^{l/s-1}(\operatorname{ln}N)^{(s-l)D},&l<s,\\ CN^{l/s-1}(\operatorname{ln}N)^{(l-s)\delta},&l>s,\\ C\operatorname{ln}\operatorname{ln}N,&l=s,\\ \end{array}\right.$
@@
(i) Suppose that $\operatorname{Sep}(n_{1},\dots,n_{k})=k.$ Then we use the multiple mixing bound of Lemma 2.4(ii). We take $K$ in the definition of separation condition so large that the error term in Lemma 2.4(ii) is $O(N^{-k}).$ There are two further subcases.
(a) $\min_{j}m_{j}>1.$ Using the second estimate in (3.30) we see that
@@
while the error term is much smaller. Summing over all $n_{1},\dots,n_{k}$ we obtain this case gives the leading contribution of $CN^{m/s}(\operatorname{ln}N)^{m\delta}.$
(b) $\min_{j}m_{j}=1.$ If $1<s<2,$ then
$\prod_{j}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)=0,$
so each term is $O(N^{-k})$ and the total contribution of all terms is $O(1).$
If $s=1,$
@@
$\leq$ $C\left(\prod_{m_{j}\geq 2}N^{m_{j}-1}\left(\operatorname{ln}N\right)^{(m_{j}-1)\delta}\right)\left(\operatorname{ln}\operatorname{ln}N\right)^{m-k}\leq CN^{m-k}\left(\operatorname{ln}N\right)^{(m-k)\delta}\left(\operatorname{ln}\operatorname{ln}N\right)^{m}$
@@
(ii) Suppose that $\operatorname{Sep}(n_{1},\dots,n_{k})=\bar{k}<k.$ Given that $x$ is Diophantine, we consider the case where
@@
for some $\epsilon>0.$ Otherwise, the contribution to the sum is $0$ By (3.30),
@@
In addition using Lemma 2.4(i), we see that $E(\bar{\bar{X}}_{n_{1}}^{m_{1}}\dots\bar{\bar{X}}_{n_{k}}^{m_{k}})$ is $O\left(N^{(m/s)-\bar{k}-\eta}(\operatorname{ln}N)^{Dm}\right),$ where $\eta=(k-\bar{k})\epsilon|\operatorname{ln}\theta|.$ Since the number of such terms in (3.29) is $O\left(N^{\bar{k}}(\operatorname{ln}N)^{k-\bar{k}}\right),$ we conclude that the contribution of these non-separated terms is $O\left(N^{m/s-\eta}(\operatorname{ln}N)^{(D+1)m}\right)$, which is much smaller than the RHS of (3.28). $\square$
The proof of (3.23) is similar to the proof of (3.14). Namely, as in Section 3.3, it suffices to have the following analogue of (3.20):
@@
for some $C>0.$ Split ${\bar{{\bar{S}}}}_{N-n}={\bar{{\bar{S}}}}_{K\operatorname{ln}N}+\left({\bar{{\bar{S}}}}_{N-n}-{\bar{{\bar{S}}}}_{K\operatorname{ln}N}\right).$ The first term is estimated by (3.27). For the second term arguing in the same way as in the proof of (3.20) but replacing (3.18) by (3.28), we obtain
@@
Therefore,
(3.32) $\mu\left(z\in I_{n}:\left|{\bar{{\bar{S}}}}_{N-n}(z)-{\bar{{\bar{S}}}}_{K\operatorname{ln}N}(z)\right|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)$
$\leq\frac{16C\mu(I_{n})}{{\varepsilon}^{2}}(\operatorname{ln}N)^{-m(\alpha-\delta)}.$
Then we get (3.31) by combining (3.27) with (3.32).
@@ Appendix A Exponential mixing of all orders for expanding maps
We prove part(i) by induction. When $q=2$ the statement reduces to (2.4). For $q>2$ we apply (2.4) with $\psi_{1}=\varphi_{1},$ $\psi_{2}=\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}-k_{2}})$ and obtain
@@
By induction
$\|\psi_{2}\|_{L^{1}}\leq\left(\prod_{j=2}^{q-1}\left(||\varphi_{j}||_{L^{1}}+\theta^{k_{j}-k_{j-1}}\|\varphi_{j}\|_{BV}\right)\right)\|\varphi_{q}\|_{L^{1}}.$
@@
Proceeding in a similar way, we can prove that
@@
Now we turn to the proof of part (ii). When $q=2$ the statement also reduces to (2.4). Applying (2.4) with $\psi_{1}=\varphi_{1},$ $\psi_{2}=\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}-k_{2}}),$ we get
@@
where the last inequality is due to part (i) of the lemma, which we have already proved. Applying inductive estimate to ${\mathbb{E}}\left(\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}})\right)$ we obtain (2.6). $\square$
@@ Appendix B Good targets for expanding maps
We prove for $\Omega_{\rho_{N},j},$ and the proof for $\Omega_{\rho_{N}}$ follows in the same approach. Denote $B(x,r_{1},r_{2})=\{y\in\mathbb{S}:r_{1}\leq|y-x|\leq r_{2}\},$ and $\tilde{A}^{+}_{\rho_{N},j},$ $\tilde{A}^{-}_{\rho_{N},j}$ to be the characteristic functions of
@@
respectively. Then
@@
$\tilde{A}^{-}_{\rho_{N},j}\leq{\mathbf{1}}_{\Omega_{\rho_{N},j}}\leq\tilde{A}^{+}_{\rho_{N},j},\quad E\tilde{A}^{+}_{\rho_{N},j}-E\tilde{A}^{-}_{\rho_{N},j}\leq C\rho_{N}^{-(s+1)}\text{ and }\|\tilde{A}^{*}_{\rho_{N},j}\|_{BV}\leq 2\mu(\Omega_{\rho_{N},j})+4,$
[…]
@@
First we prove $(M1)_{r}.$ If $0<k_{1}<k_{2}<\dots<k_{r}\leq N$ are such that $k_{i+1}-k_{i}\geq K\operatorname{ln}N,$ where $K$ is a sufficiently large constant, and each $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r},$
@@
$\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right)\leq{\mathbb{E}}\left(\prod_{j=1}^{r}\tilde{A}^{+}_{\rho_{N},i_{j}}\circ f^{k_{j}}\right)\leq$
@@
$\left(\prod_{j=1}^{r-1}\left({\mathbb{E}}\tilde{A}^{+}_{\rho_{N},i_{j}}+C\theta^{K\operatorname{ln}N}(2{\mathbb{E}}\tilde{A}^{+}_{\rho_{N},i_{j}}+4)\right)\right)({\mathbb{E}}\tilde{A}^{+}_{\rho_{N},i_{r}})\leq\left(\prod_{j=1}^{r}\sigma_{\rho_{N},i_{j}}\right)(1+{\varepsilon}_{N})$
for ${\varepsilon}_{N}\rightarrow 0,$ which yields the RHS of $(M1)_{r}.$ The LHS is proved similarly replacing $\tilde{A}^{+}_{\rho_{N},i_{j}}$ by $\tilde{A}^{-}_{\rho_{N},i_{j}}$.
Second we show $(M2)_{r}$ assuming that $x$ is Diophanitine. Let $\operatorname{Sep}(k_{1},\dots k_{r})=m<r.$ Since $x$ is Diophantine, there exists $\epsilon>0$ such that $B(x,c\rho_{N}^{-s})\cap f^{-k}B(x,c\rho_{N}^{-s})=\emptyset$ for $k\leq\epsilon\operatorname{ln}N$ and $c=2a^{s}/c_{j}^{s}.$ Thus we assume that $k_{j+1}-k_{j}\geq\epsilon\operatorname{ln}N$ for all $j.$ Then
$\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right)$$\leq$ $C\left(\prod_{k_{j+1}-k_{j}>K\operatorname{ln}N}\sigma_{\rho_{N},i_{j}}\right)(\theta^{\epsilon\operatorname{ln}N})^{r-m}(1+{\varepsilon}_{N})\leq{\frac{C\sigma_{{\rho}_{N}}^{m}}{(\operatorname{ln}N)^{100r}}}.$
@@
Third we prove $(M3)_{r}.$ If $k_{1}<\dots<k_{r}<l_{1}<\dots<l_{r}$ are such that $2^{i}<k_{\alpha}\leq 2^{i+1},2^{j}<l_{\beta}\leq 2^{j+1}$ for $1\leq\alpha,\beta\leq r$, then for each $(i_{1},i_{2},\dots,i_{r}),\,(j_{1},j_{2}\dots j_{r})\in\{1,\dots,p\}^{r},$
@@
$\mu\left(\left[\bigcap_{\alpha=1}^{r}\Omega^{k_{\alpha}}_{{\rho}_{2^{i}},i_{\alpha}}\right]\bigcap\left[\bigcap_{\beta=1}^{r}\Omega^{l_{\beta}}_{{\rho}_{2^{j}},j_{\beta}}\right]\right)\leq{\mathbb{E}}\left(\left(\prod_{\alpha=1}^{r}\tilde{A}^{+}_{\rho_{2^{i}},i_{\alpha}}\circ f^{k_{\alpha}}\right)\left(\prod_{\beta=1}^{r}\tilde{A}^{+}_{\rho_{2^{j}},j_{\beta}}\circ f^{k_{\beta}}\right)\right)$
@@
$\leq\left(\prod_{\alpha=1}^{r}\left({\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{i}},i_{\alpha}}+C\theta^{Ki}(2{\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{i}},i_{\alpha}}+4)\right)\right)\left(\prod_{\beta=1}^{r-1}\left({\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{j}},j_{\beta}}+C\theta^{Kj}(2{\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{j}},j_{\beta}}+4)\right)\right)({\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{j}},j_{r}})$
$\leq\left(\prod_{\alpha=1}^{r}\sigma_{\rho_{2^{i}},i_{\alpha}}\right)\left(\prod_{\beta=1}^{r}\sigma_{\rho_{2^{j}},j_{\beta}}\right)(1+{\varepsilon}_{i}).$
Finally we prove $(M4)_{*}$ assuming that $x$ is non-periodic. The proof of $(M4)_{I_{1},\dots I_{r},i_{1},\dots,i_{r}}$ is similar and will be left to the reader. We denote by

### v3

@@ 3.1. Splitting the sum
3.3. Controlling small values of $X$ for $1\leq s<2$
We now prove (3.3) for $1\leq s<2.$
Given $y\in\mathbb{S},$ let $C_{n}(y)$ be the continuity component of $\phi 1_{\phi\leq N^{1/s}(\operatorname{ln}N)^{-D}}\circ f^{k}$ for all $k\leq n,$ containing $y.$ Then $f^{n}|_{C_{n}(y)}$ be a diffeomorphism. Define $I_{n}(y)=f^{n}C_{n}(y).$ We say that $C_{n}(y)$ is short if the length of $I_{n}(y)$ is less than $N^{-1}(\operatorname{ln}N)^{-100},$ and long otherwise.
Fix $\varepsilon>0.$ We say that $C_{n}$ is extremal if there exists $z\in C_{n}$ such that
$\left|{\bar{S}}_{n}(z)\right|\geq{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}$
and for each $k<n,$ $\left|{\bar{S}}_{k}(\cdot)\right|<{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}$ on $C_{n}.$ Note that different extremal components are disjoint. Moreover, we have
(3.13) $\left|{\bar{S}}_{n}(y)\right|\geq\frac{3}{4}{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}\quad\text{for all }y\text{ in }C_{n}.$
@@ 3.3. Controlling small values of $X$ for $1\leq s<2$
$\left|{\bar{S}}_{n}(y)-{\bar{S}}_{n}(z)\right|\leq\sum_{k=1}^{n-K\operatorname{ln}N}\left|{\bar{X}}_{k}(y)-{\bar{X}}_{k}(z)\right|+\left|{\bar{S}}_{K\operatorname{ln}N}(f^{n-K\operatorname{ln}N}y)\right|+\left|{\bar{S}}_{K\operatorname{ln}N}(f^{n-K\operatorname{ln}N}z)\right|.$
The last two terms are smaller than $2KN^{1/s}(\operatorname{ln}N)^{1-D}$ due to the cutoff in the definition of ${\bar{S}}_{n}.$ To estimate the first term, observe that since $d(f^{n}y,f^{n}z)\leq 1$ and $f$ is expanding, for $k\leq n-K\operatorname{ln}N,$ we have $d(f^{k}y,f^{k}z)\leq\theta^{K\operatorname{ln}N}.$ Since $\phi{\mathbf{1}}_{\{\phi<N^{1/s}(\operatorname{ln}N)^{-D}\}}$ is Lipschitz on its continuity components with a Lipschitz constant less than $CN^{1+1/s}(\operatorname{ln}N)^{-(1+s)D}$, it follows that if $K$ is sufficiently large, then
$\left|{\bar{X}}_{k}(y)-{\bar{X}}_{k}(z)\right|\leq N^{-100},$
which proves (3.13).
@@
If $D$ is is sufficiently large, then $\mu({\mathcal{E}}_{N})\leq{C}{(\operatorname{ln}N)^{-2}}.$
Lemma 3.2 implies (3.3). In fact, combining this lemma with the classical Borel-Cantelli lemma we conclude that with probability $1$, for sufficiently large $k$, we have
$\left|{\bar{S}}_{n}\right|<{\varepsilon}N_{k}^{1/s}(\operatorname{ln}N_{k})^{\alpha}\,\,\,\text{for}\,\,\,n\leq N_{k},\,\,\,N_{k}=2^{k}.$
Since $\varepsilon$} is arbitrary, (3.3) follows.
@@
To prove (3.15) we use Chebyshev’s inequality. By (2.1) we have
@@
Since $D(2-s)+2\alpha>3$ due to (3.2), (3.15) follows by Chebyshev’s inequality.
The proof of (3.14) follows similar ideas. Namely, in view of (3.13), it suffices to show that there exists $\tilde{C}<1$ such that
(3.19) $\mu\left(y\in C_{n}:\left|\sum_{k=n+1}^{N}{\bar{X}}_{k}(y)\right|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\tilde{C}\mu(C_{n}).$
@@
(3.20) $\mu\left(z\in I_{n}:|{\bar{S}}_{N-n}(z)|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\tilde{C}\mu(I_{n}).$
@@
Combining (3.21) and (3.22) we obtain (3.14). ∎
3.4. Controlling intermediate values of $X$ for $1\leq s<2$
Here we prove (3.4) in case $1\leq s<2.$
@@ 3.4. Controlling intermediate values of $X$ for $1\leq s<2$
(3.25) $\sharp\left\{k:\,1\leq k\leq K\operatorname{ln}N,\,\,{\bar{{\bar{X}}}}_{k}(y)\geq N^{1/s}(\operatorname{ln}N)^{-D}\right\}\leq(\operatorname{ln}N)^{(\alpha-\delta)/2}$
@@
Equation (3.24) follows from the Markov’s inequality and the following estimate.
@@
Suppose $x$ is Diophantine. For each positive integer $m,$ we have
@@
Note that by (2.1),
[…]
@@
(i) Suppose that $\operatorname{Sep}_{N}(n_{1},\dots,n_{k})=k.$ In this case, we apply the multiple mixing bound from Lemma 2.4(ii). We choose $K$ in the definition of separation condition large enough so that the error term in Lemma 2.4(ii) is $O(N^{-k}).$ This leads to two further subcases.
(a) $\min_{j}m_{j}>1.$ Using the second estimate in (3.30), we find that
@@
while the error term is much smaller. Summing over all $n_{1},\dots,n_{k},$ we see that this case gives the leading contribution of $CN^{m/s}(\operatorname{ln}N)^{m\delta}.$
(b) $\min_{j}m_{j}=1.$ If $1<s<2,$ then $\prod_{j}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)=0,$ so each term is $O(N^{-k}),$ and the total contribution of all terms is $O(1).$
If $s=1,$ we have
@@
This leads to
$\prod_{j}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)\leq C\left(\prod_{m_{j}\geq 2}N^{m_{j}-1}\left(\operatorname{ln}N\right)^{(m_{j}-1)\delta}\right)\left(\operatorname{ln}\operatorname{ln}N\right)^{m-k}\leq CN^{m-k}\left(\operatorname{ln}N\right)^{(m-k)\delta}\left(\operatorname{ln}\operatorname{ln}N\right)^{m}$
@@
(ii) Suppose that $\operatorname{Sep}_{N}(n_{1},\dots,n_{k})=\bar{k}<k.$ Given that $x$ is Diophantine, we consider the case where
@@
for some $\epsilon>0.$ Otherwise, the contribution to the sum is $0.$ By (3.30),
@@
Additionally, using Lemma 2.4(i), we find that
$E(\bar{\bar{X}}_{n_{1}}^{m_{1}}\dots\bar{\bar{X}}_{n_{k}}^{m_{k}})=O\left(N^{(m/s)-\bar{k}-\eta}(\operatorname{ln}N)^{Dm}\right),$
where $\eta=(k-\bar{k})\epsilon|\operatorname{ln}\theta|.$ Since the number of such terms in (3.29) is $O\left(N^{\bar{k}}(\operatorname{ln}N)^{k-\bar{k}}\right),$ we conclude that the contribution of these non-separated terms is $O\left(N^{m/s-\eta}(\operatorname{ln}N)^{(D+1)m}\right),$ which is much smaller than the right-hand side of (3.28). ∎
The proof of (3.23) is similar to the proof of (3.14). Namely, as in §3.3, it suffices to establish the following analogue of (3.20):
@@
for some constant $C>0.$ We split ${\bar{{\bar{S}}}}_{N-n}={\bar{{\bar{S}}}}_{K\operatorname{ln}N}+\left({\bar{{\bar{S}}}}_{N-n}-{\bar{{\bar{S}}}}_{K\operatorname{ln}N}\right).$ The first term is estimated using (3.27). For the second term, we proceed as in the proof of (3.20), but replace (3.18) with (3.28). This yields
@@
Therefore, we have
(3.32) $\mu\left(z\in I_{n}:\left|{\bar{{\bar{S}}}}_{N-n}(z)-{\bar{{\bar{S}}}}_{K\operatorname{ln}N}(z)\right|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\frac{16C\mu(I_{n})}{{\varepsilon}^{2}}(\operatorname{ln}N)^{-m(\alpha-\delta)}.$
Finally, combining (3.27) with (3.32) yields (3.31).
@@ Appendix A Exponential mixing of all orders for expanding maps
We prove part (i) by induction. When $q=2$ the statement reduces to (2.7). For $q>2$ we apply (2.7) with $\psi_{1}=\varphi_{1},$ $\psi_{2}=\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}-k_{2}}),$ and obtain
@@
By the induction hypothesis, we have
[…]
@@
Proceeding in a similar way, we can also prove that
@@
Now we turn to the proof of part (ii). When $q=2,$ the statement reduces to (2.7). Applying (2.7) with $\psi_{1}=\varphi_{1}$ and $\psi_{2}=\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}-k_{2}}),$ we obtain
@@
where the last inequality follows from part (i) of the lemma, which we have already proven. Applying the inductive estimate to ${\mathbb{E}}\left(\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}})\right)$ we obtain (2.9). ∎
@@ Appendix B Good targets for expanding maps
We use the notation $\Omega_{\rho_{N}}$ and $\Omega_{\rho_{N},j}$ as defined in (2.4) and (2.5), respectively.
We define the annular set $B(x,r_{1},r_{2})=\{y\in\mathbb{S}:r_{1}\leq|y-x|\leq r_{2}\},$ and let $\tilde{A}^{+}_{\rho_{N},j}$, $\tilde{A}^{-}_{\rho_{N},j}$ be the characteristic functions of the sets
@@
respectively, where $L=\max_{x\in\mathbb{S}}\left|\psi(x)\right|.$ Then
@@
$\tilde{A}^{-}_{\rho_{N},j}\leq{\mathbf{1}}_{\Omega_{\rho_{N},j}}\leq\tilde{A}^{+}_{\rho_{N},j},\quad{\mathbb{E}}\tilde{A}^{+}_{\rho_{N},j}-{\mathbb{E}}\tilde{A}^{-}_{\rho_{N},j}\leq C\rho_{N}^{-(s+1)}\text{ and }\left\|\tilde{A}^{\pm}_{\rho_{N},j}\right\|_{BV}\leq 2\mu(\Omega_{\rho_{N},j})+4.$
We now verify the conditions $(M1)_{r},$ $(M2)_{r},$ $(M3)_{r},$ $(M4)_{*}$ and $(M4)_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}$ using Lemma 2.4.
@@
(a) We prove $(M1)_{r}$ and $(M3)_{r}$ for $\Omega_{\rho_{N},j},$ and the proof for $\Omega_{\rho_{N}}$ follows in the same approach. Suppose $0<k_{1}<k_{2}<\dots<k_{r}\leq N$ are such that $k_{i+1}-k_{i}\geq K\operatorname{ln}N,$ where $K$ is a sufficiently large constant, and each $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r}.$ Then we have
@@
$\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right)\leq{\mathbb{E}}\left(\prod_{j=1}^{r}\tilde{A}^{+}_{\rho_{N},i_{j}}\circ f^{k_{j}}\right)$
@@
$\leq$ $\left(\prod_{j=1}^{r-1}\left({\mathbb{E}}\tilde{A}^{+}_{\rho_{N},i_{j}}+C\theta^{K\operatorname{ln}N}\left(2{\mathbb{E}}\tilde{A}^{+}_{\rho_{N},i_{j}}+4\right)\right)\right){\mathbb{E}}\tilde{A}^{+}_{\rho_{N},i_{r}}\leq\left(\prod_{j=1}^{r}\sigma_{\rho_{N},i_{j}}\right)(1+{\varepsilon}_{N})$
where ${\varepsilon}_{N}\to 0$ as $N\to\infty.$ This establishes the right-hand side of $(M1)_{r}.$ The proof for the left-hand side follows similarly, replacing $\tilde{A}^{+}_{\rho_{N},i_{j}}$ by $\tilde{A}^{-}_{\rho_{N},i_{j}}$.
@@
Next we prove $(M3)_{r}.$ Suppose $k_{1}<\dots<k_{r}<l_{1}<\dots<l_{r}$ are such that $2^{i}<k_{\alpha}\leq 2^{i+1}$ and $2^{j}<l_{\beta}\leq 2^{j+1}$ for $1\leq\alpha,\beta\leq r.$ For each pairs $(i_{1},i_{2},\dots,i_{r})$ and $(j_{1},j_{2}\dots j_{r})$ in $\{1,\dots,p\}^{r},$ we have
@@
$\mu\left(\left(\bigcap_{\alpha=1}^{r}\Omega^{k_{\alpha}}_{{\rho}_{2^{i}},i_{\alpha}}\right)\bigcap\left(\bigcap_{\beta=1}^{r}\Omega^{l_{\beta}}_{{\rho}_{2^{j}},j_{\beta}}\right)\right)\leq{\mathbb{E}}\left(\left(\prod_{\alpha=1}^{r}\tilde{A}^{+}_{\rho_{2^{i}},i_{\alpha}}\circ f^{k_{\alpha}}\right)\left(\prod_{\beta=1}^{r}\tilde{A}^{+}_{\rho_{2^{j}},j_{\beta}}\circ f^{k_{\beta}}\right)\right)$
@@
$\leq$ $\left(\prod_{\alpha=1}^{r}\left({\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{i}},i_{\alpha}}+C\theta^{Ki}\left(2{\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{i}},i_{\alpha}}+4\right)\right)\right)\left(\prod_{\beta=1}^{r-1}\left({\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{j}},j_{\beta}}+C\theta^{Kj}\left(2{\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{j}},j_{\beta}}+4\right)\right)\right){\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{j}},j_{r}}$
$\leq$ $\left(\prod_{\alpha=1}^{r}\sigma_{\rho_{2^{i}},i_{\alpha}}\right)\left(\prod_{\beta=1}^{r}\sigma_{\rho_{2^{j}},j_{\beta}}\right)(1+{\varepsilon}_{i}),$
which proves $(M3)_{r}.$
(b) We prove $(M2)_{r}$ under the assumption that $x$ is Diophantine. Let $\operatorname{Sep}_{N}(k_{1},\dots k_{r})=m<r.$ Since $x$ is Diophantine, there exists $\epsilon>0$ such that
$B\left(x,c\rho_{N}^{-s}\right)\cap f^{-k}B\left(x,c\rho_{N}^{-s}\right)=\emptyset$
for positive integer $k\leq\epsilon\operatorname{ln}N,$ where $c=2a^{s}/c_{j}^{s}.$ Therefore, we assume that $k_{j+1}-k_{j}\geq\epsilon\operatorname{ln}N$ for all $j.$ Then we obtain